An object O is kept in air in front of a thin plano convex lens of radius of curvature 10 cm. Its refractive 


index is 3/2


and the medium towards right of the plane surface is water of refractive index 4/3.


 What 


should be the distance ‘x’ of the object so that the rays become parallel finally?


Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

for refraction at plane surface,

1/x -(3/2v1)=0,v1=3x/2

at curved surface, (3/2v1)+(4/3v2)=((4/3)-(3/2))/(-10)

v2→∞,therefore (3/2v1)=1/x=1/6

x=6cm

Vishrut Jain
6 years ago
(N(air)/v)-(N(air)/x) = [(Ng Na)/R1]+[(Nw-Ng)/R2)Here, R2-----> infinity v-----> infinitySo,-1/x = (1.5 -1)/10X = -20cm